Write the degeneracy of isotropic 3 dimensional harmonic oscillator.


1
Expert's answer
2021-02-15T00:53:40-0500

For the three-dimensional isotropic harmonic oscillator the energy eigenvalues are E = (n + 3/2)ħω, with n = n1 + n2 + n3, where n1, n2, n3 are the numbers of quanta associated with oscillations along the Cartesian axes.